# SVDQuant: Absorbing Outliers by Low-Rank Components for 4-Bit Diffusion Models

### v3

@@ 4.3 Nunchaku: Fusing Low-Rank and Low-Bit Branch Kernels
Although the low-rank branch introduces negligible computation in theory, running it as a separate branch incurs large latency overhead—approximately 50% of the 4-bit branch latency, as shown in Figure 6(a). This occurs because, for a small rank $r$, even though the computational cost decreases significantly, the input and output activation sizes remain unchanged, shifting the bottleneck from computation to memory access. This issue worsens, especially when the activation cannot fit into the GPU L2 cache. For example, in the diffusion transformer block, the up projection in the low-rank branch for QKV projection is much slower since its output exceeds the available L2 cache, resulting in the extra DRAM load and store operations. Fortunately, the down projection ${\bm{L}}_{1}$ in the low-rank branch shares the same input as the quantization kernel in the low-bit branch, while the up projection ${\bm{L}}_{2}$ shares the same output as the 4-bit computation kernel, as illustrated in Figure 6(b). By fusing the down projection with the quantization kernel and the up projection with the 4-bit computation kernel, the low-rank branch can share the activations with the low-bit branch, eliminating the extra memory access and halving the number of kernel calls. As a result, our low-rank branch adds only 5$\sim$10latency, making it nearly cost-free.
@@ 5.1 Setups
Models. We benchmark our methods using FLUX.1 (Black-Forest-Labs, 2024), PixArt-$\Sigma$ (Chen et al., 2024a), SANA (Xie et al., 2025), Stable Diffusion XL (SDXL) (Podell et al., 2024) and SDXL-Turbo (Sauer et al., 2024), including both the UNet (Ronneberger et al., 2015; Ho et al., 2020) and DiT (Peebles & Xie, 2023) backbones. See Appendix B for more details.
Datasets. Following previous works (Li et al., 2023a; Zhao et al., 2024c; b), we randomly sample the prompts in COCO Captions 2024 (Chen et al., 2015) for calibration. To evaluate the generalization capability of our method, we sample 5K prompts from the MJHQ-30K (Li et al., 2024a) and the summarized Densely Captioned Images (sDCI) (Urbanek et al., 2024) for benchmarking. See Appendix C for more details.
@@
  • 4-bit NormalFloat (NF4) is an information-theoretically optimal 4-bit data type for weight-only quantization (Dettmers et al., 2023), which assumes that weights follow a normal distribution. We use the community-quantized NF4 FLUX.1 models (Lllyasviel, 2024) as the baselines.
  • ViDiT-Q (Zhao et al., 2024b) uses per-token quantization and smoothing (Xiao et al., 2023) to alleviate the outliers across different batches and tokens and achieves lossless 8-bit quantization on PixArt-$\Sigma$.
  • MixDQ (Zhao et al., 2024c) identifies the outliers in the begin-of-sentence token of text embedding and protects them with 16-bit pre-computation. This method enables up to W4A8 quantization with negligible performance degradation on SDXL-Turbo.
@@
Metrics. Following previous works (Li et al., 2022; 2024b), we evaluate image quality and image similarity with respect to the 16-bit models’ results. For image quality assessment, we use Fréchet Inception Distance (FID, lower is better) to measure the distribution distance between the generated images and the ground-truth images (Heusel et al., 2017; Parmar et al., 2022). Besides, we employ Image Reward (higher is better) to approximate the human rating of the generated images (Xu et al., 2024a). We use LPIPS (lower is better) to measure the perceptual similarity (Zhang et al., 2018) and Peak Signal Noise Ratio (PSNR, higher is better) to measure the numerical similarity of the images from the 16-bit models. Please refer to our Appendix E.1 for more metrics (CLIP IQA (Wang et al., 2023b), CLIP Score (Hessel et al., 2021) and SSIM§§§https://en.wikipedia.org/wiki/Structural_similarity_index_measure).
@@
FLUX.1 -schnell (4 Steps) BF16 19.2 0.938 20.8 0.932
@@
PixArt-$\Sigma$ (20 Steps) FP16 16.6 0.944 24.8 0.966
@@
UNet SDXL -Turbo (4 Steps) FP16 24.3 0.845 24.7 0.705
@@ 5.2 Results
Integrate with LoRA. Previous quantization methods require fusing the LoRA branches and re-quantizing the model when integrating LoRAs. In contrast, our Nunchaku eliminates redundant memory access, allowing adding a separate LoRA branch. In practice, we can fuse the LoRA branch into our low-rank branch by slightly increasing the rank, further enhancing efficiency. In Figure 9, we exhibit some visual examples of applying LoRAs of five different styles (Realism, Ghibsky Illustration, Anime, Children Sketch, and Yarn Art) to our INT4 FLUX.1-dev model. Our INT4 model successfully adapts to each style while preserving the image quality of the 16-bit version. For more visual examples, see Appendix E.2. For FLUX.1-schnell, we further support LoRAs from one-step conditional model pix2pix-turbo (Parmar et al., 2024), enabling additional controls like sketch. An interactive demo is available here.
Ablation study. In Figure 10, we present several ablation studies of SVDQuant on PixArt-$\Sigma$. First, both SVD-only and naïve quantization perform poorly in the 4-bit setting, resulting in severe quality degradation. While applying smoothing to the quantization slightly improves image quality compared to naïve quantization, the results remain unsatisfactory. LoRC (Yao et al., 2024) introduces a low-rank branch to compensate for quantization errors, but this approach is suboptimal, as quantization errors exhibit a well-spread distribution of singular values. Consequently, low-rank compensation fails to effectively mitigate these errors, as discussed in Section 4.2. In contrast, we first decompose the weights and quantize only the residual. As demonstrated in Figure 5, the first several singular values are significantly larger than the rest, allowing us to shift them to the low-rank branch, effectively reducing weight magnitude. Finally, smoothing consolidates the outliers, enabling the low-rank branch to absorb outliers from the activations and substantially improving image quality.
@@ Appendix A Proofs
where $\sigma$ is the std deviation of the normal distribution. Equation 12 comes from the maximal inequality of Gaussian variables (Lemma 2.3 in Massart (2007)).
@@ Appendix B Benchmark Models
  • FLUX.1 (Black-Forest-Labs, 2024) is the SoTA open-sourced DiT-based diffusion model. It consists of 19 joint attention blocks (Esser et al., 2024) and 38 parallel attention blocks (Dehghani et al., 2023), totaling 12B parameters. We evaluate both the 50-step guidance-distilled (FLUX.1-dev) and 4-step timestep-distilled (FLUX.1-schnell) variants.
  • PixArt-$\Sigma$ (Chen et al., 2024a) is another DiT-based model. Instead of using joint attention, it stacks 28 attention blocks composed of self-attention, cross-attention, and feed-forward layers, amounting to 600M parameters. We evaluate it on the default 20-step setting.
  • SANA (Xie et al., 2025) is a 1.6B DiT model. It utilizes a 32× compression autoencoder (Chen et al., 2025) and replaces Softmax attention with linear attention to accelerate image generation.
  • Stable Diffusion XL (SDXL) is a widely-used UNet-based model with 2.6B parameters (Podell et al., 2024). It predicts noise with three resolution scales. The highest-resolution stage is processed entirely by ResBlocks (He et al., 2016), while the other two stages jointly use ResBlocks and attention layers. Like PixArt-$\Sigma$, SDXL uses cross-attention layers for text conditioning. We evaluate it in the 30-step setting, along with its 4-step distilled variant, SDXL-Turbo (Sauer et al., 2024).
@@ Appendix C Benchmark Datasets
  • MJHQ-30K (Li et al., 2024a) consists of 30K samples from Midjourney with 10 common categories, 3K samples each. We randomly select 5K prompts from this dataset to evaluate model performance on artistic image generation.
  • Densely Captioned Images (DCI) (Urbanek et al., 2024) is a dataset containing $\sim$8K images with detailed human-annotated captions, averaging over 1,000 words. For our experiments, we use the summarized version (sDCI), where captions are condensed to 77 tokens using large language models (LLMs) to accommodate diffusion models. Similarly, we randomly sample 5K prompts for realistic image generation.
@@ Appendix D Implementation Details
For the 8-bit setting, we use per-token dynamic activation quantization and per-channel weight quantization with a low-rank branch of rank 16. For the 4-bit setting, we adopt per-group symmetric quantization for both activations and weights, along with a low-rank branch of rank 32. INT4 quantization uses a group size of 64 with 16-bit scales. We use NVFP4 for FP4 quantization, which has native hardware support of group size of 16 with FP8 scales on Blackwell GPUs (NVIDIA Corporation, 2025). We use GPTQ (Frantar et al., 2023) to quantize the residual weights. For FLUX.1 models, the inputs of linear layers in adaptive normalization are kept in 16 bits (i.e., W4A16). For other models, key and value projections in the cross-attention are retained at 16 bits since their latency only covers less than 5% of total runtime.
The smoothing factor $\lambda\in\mathbb{R}^{m}$ is a per-channel vector whose $i$-th element is computed as $\lambda_{i}=\max(|{\bm{X}}_{:,i}|)^{\alpha}/\max(|{\bm{W}}_{i,:}|)^{1-\alpha}$ following SmoothQuant (Xiao et al., 2023) Here, ${\bm{X}}\in\mathbb{R}^{b\times m}$ and ${\bm{W}}\in\mathbb{R}^{m\times n}$. It is decided offline by searching for the best migration strength $\alpha$ for each layer to minimize the layer output mean squared error (MSE) after SVD on the calibration dataset.
@@ E.1 Visual Quality Results
We report extra quantitative quality results with additional metrics in Table 2. Specifically, CLIP IQA (Wang et al., 2023b) and CLIP Score (Hessel et al., 2021) assesses the image quality and text-image alignment with CLIP (Radford et al., 2021), respectively. Structural Similarity Index Measure (SSIM) is used to measure the luminance, contrast, and structure similarity of images produced by our 4-bit model against the original 16-bit model. We also visualize more qualitative comparsions in Figures 11, 12, 13, 14 and 15.
@@
FLUX.1 -schnell (4 Steps) BF16 0.938 26.6 0.932 26.2
@@
PixArt-$\Sigma$ (20 Steps) FP16 0.944 26.8 0.966 26.1
@@
UNet SDXL -Turbo (4 Steps) FP16 0.926 26.5 0.913 26.5
@@ E.3 Additional Ablation of SVDQuant
In Table 3, we provide additional quantitative ablation results of SVDQuant on the MJHQ prompt set (Li et al., 2024a). Across all models, NVFP4 outperforms INT4 due to its native support for smaller microscaling group sizes on Blackwell. SVDQuant leverages a low-rank branch to absorb outliers, significantly enhancing image quality in all settings. Additionally, it can incorporate GPTQ (Frantar et al., 2023) instead of round-to-nearest for weight quantization, further improving quality in most cases. Notably, combining SVDQuant with NVFP4 precision achieves the best results, reaching a PSNR of 21.5 on FLUX.1-dev, closely matching the image quality of the original 16-bit model. In Figure 17, we provide qualitative comparisons across different precision settings.
@@ E.4 Latency Results
In Table 4, we compare FLUX latency on a laptop-level 4090 GPU across different precisions. Compared to INT8, 4-bit quantization delivers a 1.3×1.3$\times$1.3 × speedup. However, without optimization, SVDQuant incurs an 18% overhead due to the low-rank branch. By eliminating redundant memory access, Nunchaku achieves latency comparable to naive INT4.
@@ E.6 Trade-off between Quality and Bitwidth
We evaluate LPIPS across different bitwidths for various quantization methods on PixArt-$\Sigma$ and FLUX.1-schnell using the MJHQ dataset in Figure 19, with weights and activations sharing the same bitwidth. Following the convention (Xiao et al., 2023; Lin et al., 2024; 2025; Li et al., 2023a; Zhao et al., 2024d; Dettmers et al., 2022), for bitwidths above 4, we apply per-channel quantization; for 4 or below, we use per-group quantization (group size 64). SVDQuant consistently outperforms naive quantization and SmoothQuant. Notably, on PixArt–$\Sigma$ and FLUX.1-schnell, our 4-bit results match 7-bit and 6-bit naive quantization, respectively.
Our SVDQuant can still generate images in the 3-bit settings on both PixArt-$\Sigma$ and FLUX.1-schnell, performing much better than SmoothQuant. Below this precision (e.g., W2A4 or W4A2), SVDQuant cannot produce images either since 2-bit symmetric quantization is essentially a ternary quantization. Prior work (Ma et al., 2024a; Wang et al., 2023a) has shown that ternary neural networks require quantization-aware training even for weight-only quantization to adapt the weights and activations to the low-bit distribution.
@@ Appendix F Text Prompts
a man in armor with a beard and a swordGHIBSKY style, a fisherman casting a line into a peaceful village lake surrounded by quaint cottagesgirl, neck tuft, white hair, sheep horns, blue eyes, nm22 stylesketched style, A squirrel wearing glasses and reading a tiny book under an oak treea panda playing in the snow, yarn art styleThe text prompts we use in Figure 16 are (in the rasterizing order):
A male secret agent in a tuxedo, holding a gun, standing in front of a burning buildingA handsome man in a suit, 25 years old, cool, futuristicA knight in shining armor, standing in front of a castle under siegeA knight fighting a fire-breathing dragon in front of a medieval castle, flames and smokeA male wizard with a long white beard casting a lightning spell in the middle of a stormA young woman with long flowing hair, standing on a mountain peak at dawn, overlooking a misty valleyGHIBSKY style, a cat on a windowsill gazing out at a starry night sky and distant city lightsGHIBSKY style, a quiet garden at twilight, with blooming flowers and the soft glow of lanterns lighting up the pathGHIBSKY style, a serene mountain lake with crystal-clear water, surrounded by towering pine trees and rocky cliffsGHIBSKY style, an enchanted forest at night, with glowing mushrooms and fireflies lighting up the underbrushGHIBSKY style, a peaceful beach town with colorful houses lining the shore and a calm ocean stretching out into the horizonGHIBSKY style, a cozy living room with a view of a snow-covered forest, the fireplace crackling and a blanket draped over a comfy chaira dog wearing a wizard hat, nm22 anime stylea girl looking at the stars, nm22 anime stylea fish swimming in a pond, nm22 stylea giraffe with a long scarf, nm22 stylea bird sitting on a branch, nm22 minimalist stylea girl wearing a flower crown, nm22 stylesketched style, A garden full of colorful butterflies and blooming flowers with a gentle breeze blowingsketched style, A beach scene with kids building sandcastles and seagulls flying overheadsketched style, A hot air balloon drifting peacefully over a patchwork of fields and forests belowsketched style, A sunny meadow with a girl in a flowy dress chasing butterfliessketched style, A little boy dressed as a pirate, steering a toy ship on a small streamsketched style, A small boat floating on a peaceful lake, surrounded by trees and mountainsa hot air balloon flying over mountains, yarn art stylea cat chasing a butterfly, yarn art stylea squirrel collecting acorns, yarn art stylea wizard casting a spell, yarn art stylea jellyfish floating in the ocean, yarn art stylea sea turtle swimming through a coral reef, yarn art style

### v4

@@ 4.3 Nunchaku: Fusing Low-Rank and Low-Bit Branch Kernels
[…]
@@ 5.1 Setups
Models. We benchmark our methods using FLUX.1 (flux1), PixArt-$\Sigma$ (chen2024pixart), SANA (xie2024sana), Stable Diffusion XL (SDXL) (podell2023sdxl) and SDXL-Turbo (sauer2023adversarial), including both the UNet (ronneberger2015u; ho2020denoising) and DiT (peebles2023scalable) backbones. See Appendix B for more details.
Datasets. Following previous works (li2023q; zhao2024mixdq; zhao2024vidit), we randomly sample the prompts in COCO Captions 2024 (chen2015microsoft) for calibration. To evaluate the generalization capability of our method, we sample 5K prompts from the MJHQ-30K (li2024playground) and the summarized Densely Captioned Images (sDCI) (urbanek2024picture) for benchmarking. See Appendix C for more details.
@@
  • 4-bit NormalFloat (NF4) is an information-theoretically optimal 4-bit data type for weight-only quantization (dettmers2023qlora), which assumes that weights follow a normal distribution. We use the community-quantized NF4 FLUX.1 models (Lllyasviel) as the baselines.
  • ViDiT-Q (zhao2024vidit) uses per-token quantization and smoothing (xiao2023smoothquant) to alleviate the outliers across different batches and tokens and achieves lossless 8-bit quantization on PixArt-$\Sigma$.
  • MixDQ (zhao2024mixdq) identifies the outliers in the begin-of-sentence token of text embedding and protects them with 16-bit pre-computation. This method enables up to W4A8 quantization with negligible performance degradation on SDXL-Turbo.
@@
Metrics. Following previous works (li2022efficient; li2024distrifusion), we evaluate image quality and image similarity with respect to the 16-bit models’ results. For image quality assessment, we use Fréchet Inception Distance (FID, lower is better) to measure the distribution distance between the generated images and the ground-truth images (heusel2017gans; parmar2021cleanfid). Besides, we employ Image Reward (higher is better) to approximate the human rating of the generated images (xu2024imagereward). We use LPIPS (lower is better) to measure the perceptual similarity (zhang2018perceptual) and Peak Signal Noise Ratio (PSNR, higher is better) to measure the numerical similarity of the images from the 16-bit models. Please refer to our Appendix E.1 for more metrics (CLIP IQA (wang2022exploring), CLIP Score (hessel2021clipscore) and SSIM§§§https://en.wikipedia.org/wiki/Structural_similarity_index_measure).
@@
FLUX.1
-schnell
(4 Steps)
BF16 19.2 0.938 20.8 0.932
@@
PixArt-$\Sigma$
(20 Steps)
FP16 16.6 0.944 24.8 0.966
@@
UNet
SDXL
-Turbo
(4 Steps)
FP16 24.3 0.845 24.7 0.705
@@ 5.2 Results
Integrate with LoRA. Previous quantization methods require fusing the LoRA branches and re-quantizing the model when integrating LoRAs. In contrast, our Nunchaku eliminates redundant memory access, allowing adding a separate LoRA branch. In practice, we can fuse the LoRA branch into our low-rank branch by slightly increasing the rank, further enhancing efficiency. In Figure 9, we exhibit some visual examples of applying LoRAs of five different styles (Realism, Ghibsky Illustration, Anime, Children Sketch, and Yarn Art) to our INT4 FLUX.1-dev model. Our INT4 model successfully adapts to each style while preserving the image quality of the 16-bit version. For more visual examples, see Appendix E.2. For FLUX.1-schnell, we further support LoRAs from one-step conditional model pix2pix-turbo (parmar2024one), enabling additional controls like sketch. An interactive demo is available here.
Ablation study. In Figure 10, we present several ablation studies of SVDQuant on PixArt-$\Sigma$. First, both SVD-only and naïve quantization perform poorly in the 4-bit setting, resulting in severe quality degradation. While applying smoothing to the quantization slightly improves image quality compared to naïve quantization, the results remain unsatisfactory. LoRC (yao2023zeroquant) introduces a low-rank branch to compensate for quantization errors, but this approach is suboptimal, as quantization errors exhibit a well-spread distribution of singular values. Consequently, low-rank compensation fails to effectively mitigate these errors, as discussed in Section 4.2. In contrast, we first decompose the weights and quantize only the residual. As demonstrated in Figure 5, the first several singular values are significantly larger than the rest, allowing us to shift them to the low-rank branch, effectively reducing weight magnitude. Finally, smoothing consolidates the outliers, enabling the low-rank branch to absorb outliers from the activations and substantially improving image quality.
@@ Appendix A Proofs
where $\sigma$ is the std deviation of the normal distribution. Equation 12 comes from the maximal inequality of Gaussian variables (Lemma 2.3 in massart2007concentration).
@@ Appendix B Benchmark Models
  • FLUX.1 (flux1) is the SoTA open-sourced DiT-based diffusion model. It consists of 19 joint attention blocks (esser2024scaling) and 38 parallel attention blocks (dehghani2023scaling), totaling 12B parameters. We evaluate both the 50-step guidance-distilled (FLUX.1-dev) and 4-step timestep-distilled (FLUX.1-schnell) variants.
  • PixArt-$\Sigma$ (chen2024pixart) is another DiT-based model. Instead of using joint attention, it stacks 28 attention blocks composed of self-attention, cross-attention, and feed-forward layers, amounting to 600M parameters. We evaluate it on the default 20-step setting.
  • SANA (xie2024sana) is a 1.6B DiT model. It utilizes a 32× compression autoencoder (chen2024deep) and replaces Softmax attention with linear attention to accelerate image generation.
  • Stable Diffusion XL (SDXL) is a widely-used UNet-based model with 2.6B parameters (podell2023sdxl). It predicts noise with three resolution scales. The highest-resolution stage is processed entirely by ResBlocks (he2016deep), while the other two stages jointly use ResBlocks and attention layers. Like PixArt-$\Sigma$, SDXL uses cross-attention layers for text conditioning. We evaluate it in the 30-step setting, along with its 4-step distilled variant, SDXL-Turbo (sauer2023adversarial).
@@ Appendix C Benchmark Datasets
  • MJHQ-30K (li2024playground) consists of 30K samples from Midjourney with 10 common categories, 3K samples each. We randomly select 5K prompts from this dataset to evaluate model performance on artistic image generation.
  • Densely Captioned Images (DCI) (urbanek2024picture) is a dataset containing $\sim$8K images with detailed human-annotated captions, averaging over 1,000 words. For our experiments, we use the summarized version (sDCI), where captions are condensed to 77 tokens using large language models (LLMs) to accommodate diffusion models. Similarly, we randomly sample 5K prompts for realistic image generation.
@@ Appendix D Implementation Details
For the 8-bit setting, we use per-token dynamic activation quantization and per-channel weight quantization with a low-rank branch of rank 16. For the 4-bit setting, we adopt per-group symmetric quantization for both activations and weights, along with a low-rank branch of rank 32. INT4 quantization uses a group size of 64 with 16-bit scales. We use NVFP4 for FP4 quantization, which has native hardware support of group size of 16 with FP8 scales on Blackwell GPUs (nvidia2025blockscaling). We use GPTQ (frantar-gptq) to quantize the residual weights. For FLUX.1 models, the inputs of linear layers in adaptive normalization are kept in 16 bits (i.e., W4A16). For other models, key and value projections in the cross-attention are retained at 16 bits since their latency only covers less than 5% of total runtime.
The smoothing factor $\lambda\in\mathbb{R}^{m}$ is a per-channel vector whose $i$-th element is computed as $\lambda_{i}=\max(|{\bm{X}}_{:,i}|)^{\alpha}/\max(|{\bm{W}}_{i,:}|)^{1-\alpha}$ following SmoothQuant (xiao2023smoothquant) Here, ${\bm{X}}\in\mathbb{R}^{b\times m}$ and ${\bm{W}}\in\mathbb{R}^{m\times n}$. It is decided offline by searching for the best migration strength $\alpha$ for each layer to minimize the layer output mean squared error (MSE) after SVD on the calibration dataset.
@@ E.1 Visual Quality Results
We report extra quantitative quality results with additional metrics in Table 2. Specifically, CLIP IQA (wang2022exploring) and CLIP Score (hessel2021clipscore) assesses the image quality and text-image alignment with CLIP (radford2021learning), respectively. Structural Similarity Index Measure (SSIM) is used to measure the luminance, contrast, and structure similarity of images produced by our 4-bit model against the original 16-bit model. We also visualize more qualitative comparsions in Figures 11, 12, 13, 14 and 15.
@@
FLUX.1
-schnell
(4 Steps)
BF16 0.938 26.6 0.932 26.2
@@
PixArt-$\Sigma$
(20 Steps)
FP16 0.944 26.8 0.966 26.1
@@
UNet
SDXL
-Turbo
(4 Steps)
FP16 0.926 26.5 0.913 26.5
@@ E.3 Additional Ablation of SVDQuant
In Table 3, we provide additional quantitative ablation results of SVDQuant on the MJHQ prompt set (li2024playground). Across all models, NVFP4 outperforms INT4 due to its native support for smaller microscaling group sizes on Blackwell. SVDQuant leverages a low-rank branch to absorb outliers, significantly enhancing image quality in all settings. Additionally, it can incorporate GPTQ (frantar-gptq) instead of round-to-nearest for weight quantization, further improving quality in most cases. Notably, combining SVDQuant with NVFP4 precision achieves the best results, reaching a PSNR of 21.5 on FLUX.1-dev, closely matching the image quality of the original 16-bit model. In Figure 17, we provide qualitative comparisons across different precision settings.
@@ E.4 Latency Results
In Table 4, we compare FLUX latency on a laptop-level 4090 GPU across different precisions. Compared to INT8, 4-bit quantization delivers a $1.3\times$ speedup. However, without optimization, SVDQuant incurs an 18% overhead due to the low-rank branch. By eliminating redundant memory access, Nunchaku achieves latency comparable to naive INT4.
@@ E.6 Trade-off between Quality and Bitwidth
[…]
Our SVDQuant can still generate images in the 3-bit settings on both PixArt-$\Sigma$ and FLUX.1-schnell, performing much better than SmoothQuant. Below this precision (e.g., W2A4 or W4A2), SVDQuant cannot produce images either since 2-bit symmetric quantization is essentially a ternary quantization. Prior work (ma2024era; wang2023bitnet) has shown that ternary neural networks require quantization-aware training even for weight-only quantization to adapt the weights and activations to the low-bit distribution.
@@ Appendix F Text Prompts
a man in armor with a beard and a sword {internallinenumbers*} GHIBSKY style, a fisherman casting a line into a peaceful village lake surrounded by quaint cottages girl, neck tuft, white hair, sheep horns, blue eyes, nm22 style sketched style, A squirrel wearing glasses and reading a tiny book under an oak tree a panda playing in the snow, yarn art style The text prompts we use in Figure 16 are (in the rasterizing order):
A male secret agent in a tuxedo, holding a gun, standing in front of a burning building A handsome man in a suit, 25 years old, cool, futuristic A knight in shining armor, standing in front of a castle under siege A knight fighting a fire-breathing dragon in front of a medieval castle, flames and smoke A male wizard with a long white beard casting a lightning spell in the middle of a storm {internallinenumbers*} A young woman with long flowing hair, standing on a mountain peak at dawn, overlooking a misty valley {internallinenumbers*} GHIBSKY style, a cat on a windowsill gazing out at a starry night sky and distant city lights {internallinenumbers*} GHIBSKY style, a quiet garden at twilight, with blooming flowers and the soft glow of lanterns lighting up the path {internallinenumbers*} GHIBSKY style, a serene mountain lake with crystal-clear water, surrounded by towering pine trees and rocky cliffs {internallinenumbers*} GHIBSKY style, an enchanted forest at night, with glowing mushrooms and fireflies lighting up the underbrush {internallinenumbers*} GHIBSKY style, a peaceful beach town with colorful houses lining the shore and a calm ocean stretching out into the horizon {internallinenumbers*} GHIBSKY style, a cozy living room with a view of a snow-covered forest, the fireplace crackling and a blanket draped over a comfy chair a dog wearing a wizard hat, nm22 anime style a girl looking at the stars, nm22 anime style a fish swimming in a pond, nm22 style a giraffe with a long scarf, nm22 style a bird sitting on a branch, nm22 minimalist style a girl wearing a flower crown, nm22 style {internallinenumbers*} sketched style, A garden full of colorful butterflies and blooming flowers with a gentle breeze blowing sketched style, A beach scene with kids building sandcastles and seagulls flying overhead {internallinenumbers*} sketched style, A hot air balloon drifting peacefully over a patchwork of fields and forests below sketched style, A sunny meadow with a girl in a flowy dress chasing butterflies sketched style, A little boy dressed as a pirate, steering a toy ship on a small stream {internallinenumbers*} sketched style, A small boat floating on a peaceful lake, surrounded by trees and mountains a hot air balloon flying over mountains, yarn art style a cat chasing a butterfly, yarn art style a squirrel collecting acorns, yarn art style a wizard casting a spell, yarn art style a jellyfish floating in the ocean, yarn art style a sea turtle swimming through a coral reef, yarn art style